# BlockingPy: approximate nearest neighbours for blocking of records for entity resolution

### v2

@@ Abstract
Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources. The objective of entity resolution is to link records without common unique identifiers that refer to the same entity (e.g., person, company). However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity. One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex or use phonetic algorithms. However, this approach assumes that these variables are free of errors and completely observed, which is often not the case. To address this challenge, we have developed a Python package, BlockingPy, which uses blocking via modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons. In this paper, we present the design of the package, its functionalities and two case studies related to official statistics. The presented software will be useful for researchers interested in linking data from various sources.
@@
C1 Current code version v0.1.15
@@
C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, networkx, nltk, voyager, faiss-cpu, mlpack, model2vec
@@ 1 Motivation and significance
The BlockingPy package distinguishes itself by using state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data. BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
@@ 2 Existing software covering blocking procedures
Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately.
@@
The only software that provides similar functionality is the blocklib package, which is part of the Anonlink project for privacy-preserving record linkage (cf. [16, 17]). We should also mention three recent packages: vicinity [18], which provides a similar API for the ANN algorithms described in Section 3, semhash [19], which allows deduplication of records using the ANN algorithms via the vicinity and LinkTransformer [20], which uses kNN-based embedding comparisons and offers exact-match blocking. However, none of these packages provide a way to use this information for blocking of records for probabilistic record linkage and deduplication. Thus, BlockingPy provides an extension to the existing Python package ecosystem.
@@ 3.2.1 Blocking records
The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines. To achieve this, users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input. Alternatively, users can choose to represent the input data as vector embeddings.
@@ 3.3 Software architecture
ANN Algorithms. Collection of ANN implementations with a common interface for the straightforward addition of new algorithms. The parameters of each algorithm can be adjusted by providing the Controls dictionary to the control_ann parameter within the block method. An example of this is shown in Listing 1.
@@
Input Processing. Manages the data pre-processing and transformation across multiple input formats. Converts dense matrices, sparse matrices, and raw text data to sparse data frames for efficient processing. Currently, we support transformation based on $n$-grams and vector embeddings.
@@ 5 Impact
The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one dataset. For instance, we have used this approach to deduplicate the number of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022. We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [29]. The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g. Олександр, should be Ołeksandr not Aleksandr).
This software opens up new research questions regarding optimal blocking strategies in administrative data integration, particularly in contexts where language differences and transliteration complexities exist. Researchers can now investigate how different blocking approaches affect linkage quality in multilingual settings, a question that was previously difficult to study systematically without appropriate tools. The software has meaningfully improved the pursuit of quality of official statistics by allowing for more accurate population estimates through better deduplication of records. Before implementing this solution, Statistics Poland relied on user-defined rules for linking data without identifiers, which often resulted in imprecise population flow estimates due to missed matches or false positives.

### v3

@@ Abstract
Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources. The objective of entity resolution is to link records without common unique identifiers that refer to the same entity (e.g., person, company). However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity. One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex or use phonetic algorithms. However, this approach assumes that these variables are free of errors and completely observed, which is often not the case. To address this challenge, we have developed a Python package, BlockingPy, which uses blocking using modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons. The package supports both CPU and GPU execution. In this paper, we present the design of the package, its functionalities and two case studies related to official statistics. The presented software will be useful for researchers interested in linking data from various sources.
@@
C1 Current code version v0.2.8
@@
C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, igraph, nltk, voyager, faiss-cpu, faiss-gpu-cuvs, mlpack, model2vec
@@ 1 Motivation and significance
The BlockingPy package uses state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data. BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
@@ 2 Existing software covering blocking procedures
Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately. In online documentation we show that using splink with BlockingPy together results in significantly improvement in the performance metrics by capturing comparison pairs that would otherwise be missed.
@@
The only software that provides similar functionality is the blocklib package (cf. [16, 17]) and vicinity [18]. The first uses probability signature and LSH-based Lambda Fold Redundant as described in [16]. The second uses semhash [19], which allows deduplication of records using kNN-based embedding comparisons and offers exact-match blocking. In contrast, BlockingPy allows various inputs as well as ANN algorithms and graphs, focusing on providing accurate blocking for probabilistic record linkage and deduplication. In Appendix Appendix A. Numerical comparison with blocklib, we compare BlockingPy with blocklib in terms of computational efficiency, accuracy, and number of pairs from the blocking procedures. While blocklib is faster, it results in a significantly higher number of missed pairs and several orders of magnitude more pairs. We also provide comparison with the GPU version of the package.
@@ 3.1 Installation
GPU option. We also provide blockingpy-gpu with CUDA-enabled GPU execution:
1# 1) Create env
2mamba create -n blockingpy-gpu python=3.10 -y
3conda activate blockingpy-gpu
4
5# 2) Install FAISS GPU (tested cuVS nightly)
6mamba install -c pytorch/label/nightly \
7 faiss-gpu-cuvs=1.11.0=py3.10_ha3bacd1_55_cuda12.4.0_nightly -y
8
9# 3) Install BlockingPy (GPU build)
10pip install blockingpy-gpu
Notes: Import remains import blockingpy.
@@ 3.2.1 Blocking records
The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines. Scalability is achieved mainly due to the ANN libraries that we use in our implementations. Users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input. Alternatively, users can choose to represent the input data as vector embeddings.
@@
  • 6.
    Inverted File Flat (IVF), IVF with Product Quantization (IVFPQ) and CAGRA [26] from faiss (GPU) package.
@@ 3.3 Software architecture
ANN Algorithms. This is a collection of ANN implementations with a common interface for the straightforward addition of new algorithms. The parameters of each algorithm can be adjusted by providing the Controls dictionary to the control_ann parameter within the block method. An example of this is shown in Listing 1. All ANN backends share one interface across CPU and GPU.
@@
Input Processing. Manages the data pre-processing and transformation across multiple input formats. Converts dense matrices, sparse matrices, and raw text data to Numpy arrays or Scipy CSR matrices via the DataHandler class for efficient processing. Currently, we support transformation based on $n$-grams and vector embeddings.
@@ 5 Impact
The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one dataset. For instance, we have used this approach to deduplicate records of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022. We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [29]. The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g. Олександр, should be Ołeksandr not Aleksandr). Before implementing this solution, Statistics Poland relied on user-defined rules for linking data without identifiers, which often resulted in imprecise population flow estimates due to missed matches or false positives.
@@ Acknowledgements
This software was peer-reviewed and accepted by the pyOpenSci organisation (see https://github.com/pyOpenSci/software-submission/issues/232).
@@
Appendix A. Numerical comparison with blocklib
We compare BlockingPy and blocklib in three synthetic datasets ($1,500$, $15,000$, and $150,000$ records) with duplicates $500$, $5,000$, and $50,000$, respectively. The results are averaged over 10 runs. The datasets were generated with geco3 and contain fictional personal information (given name, middle name, surname, municipality, date of birth, and nationality). CPU experiments used a 6-core Intel Core i5 with 16 GB RAM and Python 3.12. GPU runs used a NVIDIA RTX 3050 (4 GB VRAM) and Python 3.10. We report time (s), recall, reduction ratio (RR), and candidate pairs (millions). The comparison is also available in the documentation along with the code and data in the project repository. Preprocessing: for BlockingPy, we concatenate all fields of interest into a single text field; for blocklib, we tune algorithm parameters to approximately match the reduction ratio of the other methods to ease comparison.
Table 4: CPU results on three datasets.
Algorithm Size Time [s] Recall $R$ Pairs (M)
P-Sig 1,500 0.051 (0.010) 0.599 (0.000) 0.996 (0.000) 0.004 (0.000)
$\lambda$-fold LSH 1,500 0.193 (0.009) 0.466 (0.033) 0.991 (0.003) 0.010 (0.003)
BlockingPy (voyager) 1,500 0.293 (0.029) 0.952 (0.006) 0.997 (0.000) 0.003 (0.000)
BlockingPy (faiss_hnsw) 1,500 0.164 (0.049) 0.960 (0.000) 0.998 (0.000) 0.003 (0.000)
BlockingPy (faiss_lsh) 1,500 0.185 (0.015) 0.955 (0.008) 0.997 (0.000) 0.003 (0.000)
P-Sig 15,000 0.388 (0.087) 0.616 (0.000) 0.996 (0.000) 0.407 (0.000)
$\lambda$-fold LSH 15,000 2.016 (0.251) 0.454 (0.028) 0.992 (0.003) 0.940 (0.337)
BlockingPy (voyager) 15,000 6.777 (0.903) 0.876 (0.005) 1.000 (0.000) 0.041 (0.001)
BlockingPy (faiss_hnsw) 15,000 12.029 (1.681) 0.913 (0.000) 1.000 (0.000) 0.031 (0.000)
BlockingPy (faiss_lsh) 15,000 1.398 (0.148) 0.900 (0.004) 1.000 (0.000) 0.033 (0.001)
P-Sig 150,000 3.444 (0.091) 0.609 (0.000) 0.996 (0.000) 40.231 (0.000)
$\lambda$-fold LSH 150,000 19.848 (0.476) 0.450 (0.026) 0.992 (0.003) 95.064 (34.252)
BlockingPy (voyager) 150,000 110.341 (3.815) 0.715 (0.005) 1.000 (0.000) 0.675 (0.021)
BlockingPy (faiss_hnsw) 150,000 250.012 (4.232) 0.832 (0.001) 1.000 (0.000) 0.375 (0.001)
BlockingPy (faiss_lsh) 150,000 56.544 (1.315) 0.818 (0.001) 1.000 (0.000) 0.400 (0.005)
Table 5: GPU results on 150,000 records dataset.
Algorithm Size Time [s] Recall $R$ Pairs (M)
BlockingPy (gpu_faiss cagra) 150,000 51.084 (0.802) 0.827 (0.000) 1.000 (0.000) 0.380 (0.000)
BlockingPy (gpu_faiss flat) 150,000 22.511 (0.508) 0.839 (0.000) 1.000 (0.000) 0.367 (0.000)
BlockingPy (gpu_faiss ivf) 150,000 71.458 (2.185) 0.801 (0.004) 1.000 (0.000) 0.487 (0.021)